Write an equation line containing the given point and parallel to the given line. express your answer in the form y=m x+b. The equation of the line is y= ____
Simply answer. (Use integers or fraction for any numbers in the expression.
(-7,2); 8x=3y+7

Answer by ewatrrr(24785)


Hi,
Write an equation line containing Pt(-7,2) and parallel to the line 8x=3y+7
*Note: The standard point-slope form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
8x=3y+7 Solving for y to put this equation into the slope-intercept form
8x - 7 = 3y
(8/3)x -(7/3) = y
Slope is (8/3) Parellel lines have equal slopes
New line:
y = (8/3)x + b Using ordered pair(-7,2) to solve for b
2 = (8/3)*(-7) + b
6/3 +(56/3) = b
62/3 = b
the equation of the line containing Pt(-7,2) and parallel to the line 8x=3y+7
y = (8/3)x + 62/3